# Improving Zero-noise Extrapolation for Quantum-gate Error Mitigation using a Noise-aware Folding Method

### v1

@@ III.1 Overview
Previous methodologies adjusted the gate count of the quantum circuit using the scale factor ($\lambda$), employing techniques such as gate-level unitary folding [30], identity gate insertion [32], and pulse stretching [29] to account for noise scaling. Although these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system. In real systems, where the error rates differ at distinct gate levels, these scaling methods cannot obtain uniform scaling factors. Such nonuniform scaling factors can impede the convergence of the extrapolation model toward a zero-noise state effectively.
@@
To tackle these problems, we propose leveraging the noise model of the target hardware to scale the noise on the quantum circuit for ZNE. Our approach regards $\lambda$ not solely as a gate-count scaling factor, but as an error-rate modifier that directly affects the quantum circuit. Quantum computers demand periodic calibration, meaning that calibration data are available for error mitigating by methods such as noise-adaptive compilation [34]. Our approach optimizes the mapping using the noise-adaptive compilation proposed by Murali et al. [34], which favors qubits with high resilience. This method maps logical qubits to physical qubits with low noise levels and minimal distance, thereby reducing the SWAP gate requirements. Employing our noise-aware folding algorithm, we then scale the approximate error rates within the transpiled circuit effectively. The following section explains the specifics of our noise-aware folding method.
III.2 Noise Accumulation
@@
Figure 2 illustrates a quantum circuit and its associated target hardware data, which are input to our proposed method. First, we transpile the quantum circuit using the noise-adaptive compilation mentioned in Section III, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates. Although noise is intentionally added to the circuit, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods. To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE research—is crucial.
@@
The transpiled circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits. Along each off-diagonal (where $i$ and $j$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates between each qubit pair (where $i$ and $j$ denote the indices of the control target qubit, respectively, with $i\neq j$). The noise of the one-qubit gates is accumulated along the diagonal matrix ($i=j$). Because the error rate of a two-qubit gate is independent of the control or target qubit, the noise in the two-qubit gates is accumulated only between qubit pairs located on the upper diagonal of the matrix.
The error rate for each pair of qubits $i$ and $j$, the error rate is aggregated only if these qubits are physically connected and if two-qubit gates exist between them. Once the error rates of the quantum circuit are aggregated into the error rate matrix, the circuit can be scaled at each scale factor.
@@
1 Function  fold_noise_aware
@@ 4       $folded\_circuit$leftarrow empty list;
       // scale the circuit in every $\lambda$ scale factor
@@ 7             $qc\_rate$leftarrow GetHighestRateIn($qc\_matrix$);
8             $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$) $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
             // Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$
@@ 11                         $cur\_rate$leftarrow $qc\_matrix$[i][j];
                         // Perform unitary folding until the adjust rate is reached
@@
                               // Get the qubit error rate using control (i) and target (j) qubit as index.
@@ 13                               $gate\_rate$leftarrow $cx\_error\_dic[i,j]$;
                               // For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$
@@ 14                               $cur\_rate$ += $gate\_rate*2$;
                               // Apply CNOT gate folding to pair i,j
@@
III.3 Noise-aware Folding
As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count. This approach involves a nuanced computation (Algorithm 1). By integrating the quantum circuit error rate matrix ($qc\_matrix$ on Line 3) with a scaling constant, we derive the $scaled\_matrix$ on Line 6, effectively delineating the desired error rate augmentation of the scaled $\lambda$ circuit.
To control the pace of noise amplification, our strategy also calculates the average disparity between the maximum error rates within the $scaled\_matrix$ and the ongoing $qc\_matrix$ (Lines 6-9 of Algorithm 1). The computed balanced error rate termed the $adjust\_rate$ on Line 9, is a regulatory mechanism that caps the rate at which errors accumulate.
Next, the error rate is amplified through the unitary folding technique, which substitutes $U$ with $U(U^{\dagger}U)$ to introduce supplementary gates. For every qubit pair within the $qc\_matrix$, this procedure iterates through the gate increments until the existing error rate of the qubit pair ($cur\_rate$) approaches or slightly surpasses the designated $adjust\_rate$ (Lines 10-17 of Algorithm 1).
The method concludes after evaluating all pairs along the upper diagonal matrix. The error rate in resultant circuit is approximately equivalent to the predefined scale $\lambda_{n}=(\lambda_{1}+(\lambda_{1}*\lambda_{n}))/2$.
Importantly, because error rates are available only for gate operations in the noise model, this study exclusively accumulates the noise in two-qubit gates. Nevertheless, our proposed methodology potentially enables seamless extension to all types of gates, provided that adequate error rate calibration data exist within the noise model.
IV Evaluation
@@ IV.1 Experimental Setup
Figure 4: Comparison of expectation values (vertical axis) obtained through using ZNE linear fit model with different folding methods on a quantum computer simulator, each with its respective noise model. (a) and (b) display results utilizing noise models from ibm_mumbai and ibm_cairo, respectively. The evaluation involves three folding methods: fold from left, fold at random, and noise-aware folding applied to the circuit depicted in Figure 3, spanning from 2 to 27 qubits (horizontal axis). The scaling factor $\lambda$ of $[1,1.5,2,2.5]$ was utilized for each folding method at every qubit count. The ‘Exact Simulation’ represents results without any applied noise model, while the ‘Unmitigated’ results showcase the circuit execution using a noise model without error mitigation. Each result is represented by a distinct symbol and color, as shown in (a).
@@
Our proposed method was validated in various full-noise models from quantum computers available on Qiskit [35]. In these evaluations, we evaluated both the setup circuit in Figure 4 and the Bernstein-Vazirani (BV) circuit as a benchmark [36]. We chose this circuit for benchmarking our proposed method for specific reasons. At a logical level, it presents a relatively straightforward circuit that alters the states of individual qubits. When errors or noise are introduced to this circuit, they are nonnegligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states. Any error during execution notably impacts the subsequent qubit within the circuit, depending on the point of its incidence. Moreover, owing to the connectivity constraints, this circuit is not easily executable on quantum computers. Accommodating these limitations in quantum computer connectivity becomes increasingly intricate with the increasing number of qubits.
@@
For each noise model, we ran the quantum circuit with different scaling parameters for comparison purposes. Along with our noise-aware folding method, we employed both from-left and at-random folding methods in [30] implemented on the Mitiq framework [37]. However. owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level. Given the complexity of our benchmark circuit (Figure 3) and the connectivity limitations of the quantum computer, we must introduce numerous SWAP gates as the number of qubits increases. To avoid this problem and ensure a fair comparison, we preemptively transpile the quantum circuit using Qiskit at the highest $optimization\_level=3$ before initiating the folding process in each method. The resulting transpiled circuit was input to each folding method, each with its own processes for extrapolation.
@@
In addition, all experimental circuits were executed on real quantum computers (see Subsection IV.3). All comparison methods, on both the simulators and real quantum computers, employed the same parameters and compilation process. The subsequent subsection delves into the evaluation results.
IV.2 Result on Simulators
@@
Our experiments were conducted on the currently available 27-qubit system, making use of the full-noise models implemented in Qiskit. Each experiment was conducted five times with different qubit counts and folding methods. Reported are the average outcomes of the five experiments for each folding method. The following graphs present the result of the linear fit extrapolation model, which demonstrated notably reliable results across all qubit counts.
@@
Figure 4 depicts the outcomes of executing the benchmark circuit Figure 3 on three full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator. As shown in panels (a) and (b) of this figure, the expectation decreased with increasing qubit counts. Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.275, 0.347, 0.350, and 0.397 in the 27-qubit circuit (Figure 3(a)).
@@
The expectation values of all execution methods, (including the unmitigated method) were high in the ibm_cairo noise model than in the ibm_mumbai (c.f. Figure 3(a) and  3(b)). This discrepancy stems from differences in the noise model parameters. Specifically, ibm_mumbai yields a significantly higher error rate than the other models.
@@
Figure 5 presents the experimental results of a 14-qubit BV circuit using the full-noise model simulation from ibm_cairo. The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but showed no consistent drop in the left and random folding methods. This result demonstrates that in our proposed method, the extrapolation model accurately extrapolates to the zero-noise value.
IV.3 Result on Real Quantum Computers
@@
Figure 6 presents the expectation values on the benchmark circuit in Figure 4, obtained through experiments on a real quantum computer ibm_algiers. The results, which mirror those of Figure 4, were averaged over five executions of each folding method. Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.979, 0.98, 0.99, and 0.979, respectively, in the 2-qubit circuit to 0.005, 0.021, 0.011, and 0.25, respectively in the 24-qubit circuit. The proposed method consistently outperformed the existing methods from the 2 to 16 qubits and performed comparably to the existing methods beyond the 16-qubit threshold. This discrepancy can be explained by the noticeably faster decline of the expectation value on scaled circuits in our proposed method than in the existing ones. At higher $\lambda$ values, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
@@
Comparing executions on real quantum computers to simulations, the overall results deteriorate significantly with increasing qubit counts on real quantum computers. This disparity arises because the noise model does not encompass all noise sources present in real quantum computers during simulations. Consequently, the results are less reliable on real quantum computers.
@@
Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [36] circuit with 14-qubit on ibm_cairo noise model simulation. The scale factor shows the drop in expectation value on different folding methods as the circuit gets nosier while $\lambda=0$ is the extrapolated expectation value where the zero-noise is assumed.
Figure 6: Expectation value results on a real quantum computer ibm_algiers with the exact same setup and scaling factor configuration from the result in Figure 4.
V Discussion

### v2

@@ 3.1 Overview
Although some of these approaches require no specific hardware noise models, they assume uniform noise scaling across the circuit, ignoring the variations in error levels among qubits within the quantum system. In real systems, where error rates differ at different logical gates for each physical qubit, these scaling methods cannot obtain uniform scaling factors. Such nonuniform scaling factors can impede the convergence of the extrapolation model toward a zero-noise state effectively.
@@
To address these problems, we propose to leverage the noise model of the target hardware to scale the noise using a noise-aware folding method in ZNE for the quantum circuit. Our approach regards $\lambda$ not solely as a gate-count scaling factor which is the main differ from fold from left method, but as an error-rate modifier threshold to scale the quantum circuit. Quantum computers require periodic calibration, which means that calibration data are readily available to the public. Although the calibration data is limited to a certain logical gates such as (one- and two-qubit gate) especially the native gate of the quantum computer, our proposed method can be used with the available data as little as a calibration error rate of a single logical two-qubit gate. Our approach optimizes the mapping using the noise-adaptive compilation proposed by Murali et al. [33], which favors qubits with high resilience. This method maps logical qubits to physical qubits with low noise levels and minimal distance, thus reducing the SWAP gate requirements. Employing our noise-aware folding algorithm, we then effectively scale the approximate error rates within the mapped circuit. The following section explains the specifics of our noise-aware folding method.
3.2 Noise Accumulation
@@
Figure 2 depicts an input quantum circuit alongside the target quantum hardware properties, encompassing the qubit connectivity topology and the calibration data for 2-qubit error rates. In this illustration, we focus on the two-qubit gate due to its higher error rate compared to a one-qubit gate. Nevertheless, our method is readily adaptable to support both one- and two-qubit gates. This visualization helps to understand how the circuit will be mapped onto the hardware and how errors may affect its execution. The error rate matrix contains a 2-qubit error rate for all of the qubits that have a connection to another qubit. First, the input quantum circuit undergoes noise-adaptive compilation, as described in Section 3, efficiently mapping the quantum circuit onto the target hardware by selecting the most robust qubits with low error rates, shown in Figure 2. In this process, logical qubits $Q_{1}$, $Q_{2}$, and $Q_{3}$ are mapped to physical qubits $1$, 00, and $2$, respectively. These specific physical qubits were chosen because their error rates for 2-qubit gates are lower compared to the rest of the qubits in the system. Although noise is intentionally added to the circuit in the next process of all of the folding methods, noise-adaptive mapping is crucial for ensuring reliable operations, minimizing qubit movement, and reducing the need for SWAP gates, which largely affect the performances of folding methods. To leverage this advantage, a deeper understanding of quantum computer noise—often overlooked in previous ZNE researches is crucial. Let’s consider a mapped circuit comprised of $d$ unitary layers. However, unlike Equation 10 where each layer $L$ represents a unitary gate, here $L$ can be either a one- or two-qubit gate. Thus, the circuit can be represented as
$U=L_{d}...L_{2}L_{1},$ (14)
where for each $j\in{1,2,...,d}$, $L_{j}$ is defined as:
$L_{j}\rightarrow\left\{\begin{array}[]{lr}L_{j}\rightarrow g_{i,k},&if~{}L_{j}\textit{is a two-qubit gate}\\ L_{j}\rightarrow g_{i},&if~{}L_{j}\textit{is a one-qubit gate}\end{array}\right.$ (15)
Here, $i$ and $k$ represent logical qubits, where $g_{i,k}$ denotes a two-qubit gate acting on the control qubit $i$ and the target qubit $k$, and $g_{i}$ represents a one-qubit gate acting solely on the target qubit $i$.
@@
The mapped circuit accumulates noises into an $n$$\times$ $n$ matrix denoted as $qc\_matrix$ in Algorithm 1, where $n$ represents the number of qubits. In the matrix, the number of rows and columns represents the physical qubit. Along each off-diagonal (where $i$ and $k$ are the rows and column indices, respectively), we aggregate the noise for two-qubit gates $g_{i,k}$ between each qubit pair (where $i$ and $k$ denote the indices of the control target qubit, respectively, with $i\neq k$) by simply summing the total error rate value using the input calibration error rate to the number of two-qubit gate $g_{i,k}$. The noise accumulation in this phase represents $\lambda=1$, indicating that the folding methods do not introduce any additional gates or errors to the circuit. The error rate for each pair of qubits $i$ and $j$, the error rate is aggregated only if these qubits are physically connected and if two-qubit gates exist between them. While along each diagonal in the $qc\_matrix$, one-qubit gate $g_{i}$ are accumulated. However in our example, in Figure 2 there are no one qubit-gate, thus making the value along the diagonal all 00.
Once the error rates of the quantum circuit are aggregated into the error rate matrix, the circuit can be scaled at each scale factor. Subsequent noise-aware folding processes for $\lambda$ are discussed in the next section.
@@
1
2Function  fold_noise_aware
3
@@ 6      $folded\_circuit$leftarrow empty list;
      /* scale the circuit in every $\lambda$ scale factor */
@@ 10            $qc\_rate$leftarrow GetHighestRateIn($qc\_matrix$);
11            $scaled\_rate$leftarrow GetHighestRateIn($scaled\_matrix$)
12            $adjust\_rate$leftarrow ($qc\_rate$ + $scaled\_rate)$ /// $2$;
            /* Scale each pair of the qubit on the upper diagonal of the $qc\_matrix$/
@@ 17                        $cur\_rate$leftarrow $qc\_matrix$[i][j];
                        /* Perform unitary folding until the adjust rate is reached */
@@
19
                              /* Get the qubit error rate using control (i) and target (j) qubit as index. */
@@ 20                               $gate\_rate$leftarrow $cx\_error\_dic[i,j]$;
                              /* For every unitary folding insertion, twice the gate $(U)$ error rate is added for $U^{\dagger}U$/
@@ 21                               $cur\_rate$ += $gate\_rate*2$;
                              /* Apply CNOT gate folding to pair i,j */
@@
3.3 Noise-aware Folding
As previously discussed, $\lambda$ in our methodology is leveraged to amplify the noise accumulation and is not directly correlated with the circuit’s gate count or depth as previously discussed on global or local unitary folding. To achieve this, we propose a noise scaling based on the error rate with the adjustment using $\lambda$ as follow:
$\epsilon_{max}=(\epsilon_{circuit}+\epsilon_{\lambda})/\gamma,$ (16)
where $\epsilon_{circuit}$ is a real number representing the highest error rate in the matrix accumulated at $\lambda=1$, as shown in Figure 2. Here, $\epsilon_{\lambda}$ is the error rate adjusted by $\lambda$, calculated as $\epsilon_{\lambda}=\epsilon_{circuit}*\lambda$. Finally, $\gamma$ is an adjustable coefficient parameter that allows control the rate of the scaling. In Figure 2 and algorithm 1, $\gamma$ is set to $2$. This computation is detailed in the algorithm 1 line 6-9.
Using $\epsilon_{max}$ at any $\lambda$, our noise-aware folding mechanism applies the unitary folding using Equation 9 until the folded gate error accumulation reaches close or equal to $\epsilon_{max}$, as detailed in algorithm 1 line 13-17.
This approach is noise-aware and significantly differs from previous methods because the rate at which unitary folding is applied to the circuit depends entirely on the error rate of the physical qubit. As illustrated in Figure 2, at $\lambda=2$, unitary folding was applied equally by adding 2 CNOT gates to pair of logical qubit $Q_{1,2}$ and $Q_{2,3}$ on the quantum circuit. However notice at $\lambda=4$, 4 CNOT gates were added to the pair of logical qubit $Q_{1,2}$ while only 2 CNOT gates were added to the pair of logical qubit $Q_{2,3}$ in the quantum circuit despite pair of $Q_{1,2}$ has more CNOT gates and a SWAP gate as well. This is because the accumulated error rate on the pair of qubit $Q_{1,2}$ and $Q_{2,3}$ are 0.063 and 0.057, respectively, with the $\epsilon_{max}=0.063$. $Q_{1,2}$ reaches the maximum threshold, while $Q_{2,3}$ reaches close to maximum and cannot add additional $U^{\dagger}U$ anymore due to it would be exceed $\epsilon_{max}$ and such condition is not allowed.
Note that this study focuses solely on applying the folding method for two-qubit gates due to their notable error rates and the availability of calibration data. However, our proposed noise-aware folding method can be expanded to accommodate all gate types, assuming sufficient error rate calibration data are available within the noise model.
4 Evaluation
@@ 4.1 Experimental Setup
Figure 4: Comparison of expectation values (vertical axis) obtained through using ZNE linear fit model with different folding methods on a quantum computer simulator, each with its respective noise model. (a) and (b) show results using noise models from ibm_mumbai and ibm_cairo, respectively. The evaluation involves three folding methods: fold from left, fold at random, and noise-aware folding applied to the circuit depicted in Figure 3, spanning from 2 to 27 qubits (horizontal axis). The scaling factor $\lambda$ of $[1,1.5,2,2.5]$ was utilized for each folding method at every qubit count. The ‘Exact Simulation’ represents results without any applied noise model, while the ‘unmitigated’ results showcase the circuit execution using a noise model without error mitigation. Each result is represented by a distinct symbol and color, as shown in (a).
@@
Our proposed method was validated in various full-noise models from quantum computers available on Qiskit [34]. In these evaluations, we evaluated both the setup circuit in Figure 3 and the Bernstein-Vazirani (BV) circuit as a benchmark [35]. We selected this circuit (see Figure 3) to assess our proposed method due to its similarity to the fully-entangled layer utilized in various NISQ applications. At a logical level, it presents a relatively straightforward circuit that alters the states of individual qubits. When errors or noise are introduced to this circuit, they are non-negligibly propagated between qubits and become significant due to their correlation and the necessity to change subsequent qubit states. Any error during execution notably impacts the subsequent qubit within the circuit, depending on the point of its incidence. Moreover, owing to the connectivity constraints, this circuit is not easily executable on quantum computers. Accommodating these limitations in quantum computer connectivity becomes increasingly complex with the increasing number of qubits.
@@
For each noise model, we ran the quantum circuit with different scaling parameters for comparison purposes. Along with our noise-aware folding method, we employed both from-left and at-random folding methods in [30] implemented on the Mitiq framework [36]. However, owing to the nature of ZNE, we must preserve the folded gates in the folded circuits, which precludes optimization such as gate cancelations before executing the quantum circuit at the compilation level. Given the complexity of our benchmark circuit (Figure 3) and the connectivity limitations of the quantum computer, we must introduce numerous SWAP gates as the number of qubits increases. To avoid this problem and ensure a fair comparison, we preemptively transpile the quantum circuit using Qiskit at the highest $optimization\_level=3$ before initiating the folding process in each method. The resulting transpiled circuit was input to each folding method, each with its own processes for extrapolation.
@@
In addition, all experimental circuits were executed on real quantum computers (see Subsection 4.3). All comparison methods, both on simulators and on real quantum computers, employed the same parameters and compilation process. The subsequent subsection delves into the evaluation results.
Figure 5: Comparison of the expectation value using different folding methods on a Bernstein-Vazirani (BV) [35] circuit with 14-qubit on ibm_cairo noise model simulation. The scale factor shows the drop in expectation value in different folding methods as the circuit gets nosier, while $\lambda=0$ is the extrapolated expectation value where zero noise is assumed.
4.2 Result on Simulators
@@
Our experiments were conducted on the 27-qubit system, making use of the full-noise models implemented in Qiskit. Each experiment was conducted five times with different qubit counts and folding methods. The results are the average of the five experiments for each folding method. The following graphs present the result of the linear fit extrapolation model, which demonstrated notably reliable results across all qubit counts.
@@
Figure 4 depicts the results of executing the benchmark circuit shown in Figure 3 on two full-noise models, namely, ibm_mumbai and ibm_cairo, using a simulator. As shown in panels (a) and (b) of this figure, the expectation decreased with increasing qubit counts. Applying the unmitigated, fold from left, random folding, and noise-aware folding to the benchmark circuit in the ibm_mumbai noise model, the expectation values respectively decreased from 0.943, 0.952, 0.949, and 0.959 in the 2-qubit circuits to 0.237, 0.292, 0.279, and 0.318 in the 27-qubit circuit (Figure 4(a)).
@@
The expectation values of all execution methods, (including the unmitigated method) were high in the ibm_cairo noise model than in the ibm_mumbai (c.f. Figure 4(a) and  4(b)). This discrepancy arises from differences in the noise model parameters. Specifically, ibm_mumbai yields a significantly higher error rate than the other models.
@@
Figure 5 presents the experimental results of a 14-qubit BV circuit using the full-noise model simulation from ibm_cairo. The expectation value reduced with increasing scaling factor $\lambda$ in our proposed method, but did not show a consistent drop in the left and random folding methods. This result demonstrates that in our proposed method, the extrapolation model accurately extrapolates to the zero-noise value.
Figure 6: Expectation value results on a real quantum computer ibmq_mumbai with the exact same setup and scaling factor configuration from the result in Figure 4.
4.3 Result on Real Quantum Computer
@@
Figure 6 presents the expectation values on the benchmark circuit shown in Figure 3, obtained through experiments on a real quantum computer ibmq_mumbai. As of preparing this manuscript, IBM has retired ibmq_mumbai; however, we captured its noise model at exact same time that we performed the experiment. You can find details regarding its properties and noise model in the appendix. The results, which mirror those of Figure 4, were averaged over five executions of each folding method. Notably, the expectation values of the unmitigated, fold from left, fold at random, and noise-aware folding methods declined from 0.949, 0.956, 0.953, and 0.981, respectively, in the 2-qubit circuit to 0.007, 0.019, 0.015, and 0.12, respectively in the 24-qubit circuit. The proposed method consistently outperformed the existing methods from the 2 to 21 qubits. This discrepancy can be explained by the noticeably faster decline of the expectation value on scaled circuits in our proposed method than in the existing ones. At higher $\lambda$ values with those of higher than 22-qubit, certain folded circuits reach a 0 expectation value, rendering the extrapolated results unreliable.
@@
Comparing executions on real quantum computers to simulations, the overall results drop significantly with increasing qubit counts on real quantum computers. This disparity arises because the noise model cannot fully represent all noise sources in real quantum computers during simulations making the results on real quantum computers less reliable.
@@
5 Discussion